# Programmable Cognitive Bias in Social Agents

### v2

@@ 3.1 Method
In the Economist condition, we modified only the profession, changing “a pharmacy shopkeeper at the Willow Market” to “a professor of economics at a nearby college,” while keeping all other details identical. The Asian Disease scenario prompt was used directly as the input.
For the role-based example (Chen et al., 2024b), we introduced two conditions: Teacher (the agent was specified as having the profession of a teacher) and Blank (no specification). We adopted the Teacher prompt from (Chen et al., 2024b) with the Problem filled using the Asian Disease scenario prompt and the Student’s Solution presented in either a positive or negative frame (see Appendix B for full prompt):
You are an experienced dialogue teacher. As a good teacher, you carefully
check the correctness of the given response based on the text. When the
solution is flawed, you should patiently teach the students how to give
better response. You must respond in the following format: Interesting:
(a score between 0 and 9) Engaging:... Problem: [description] Student’s
Solution: [solution]...
Simulated social experiment setup. We adapted the classic Asian Disease study (Tversky & Kahneman, 1981) for testing framing effect-participants tended to make more risk-averse decisions when outcomes were positively framed (e.g., “200 people will be saved”), but exhibited greater risk-seeking behavior when the same outcomes were presented as losses (e.g., “400 people will die”), into a prompt-based scenarios template. A sample scenario prompt is as follows:
You are presented with a scenario where a disease outbreak is expected to
kill 600 people unless a program is implemented. Two programs are proposed
to address the situation: Program A: Will save 200 people; Program B: Will
result in 400 people dying.
We then measured the agent’s framing preference using a five-point Likert multiple-choice scale: strong preference for the positive frame, preference for the positive frame, no preference, preference for the negative frame, and strong preference for the negative frame.
@@
Models and repetitions. We then tested the agent specifications across four models—Mistral 7B, Gemma2 9B, GPT-4o Mini, and DeepSeek-v3—using a temperature of 0.25. We intentionally used a low, more deterministic temperature of 0.25 to minimize randomness. Each example was evaluated using 15 scenario prompts, each with varying positive and negative frames, repeated 10 times, resulting in a total of 150 queries. The frequency of choice selections was recorded for analysis.
Figure 3: The example of persona-based specification (Park et al., 2023) failed to produce consistent agent behavior regarding the framing effect in the Asian Disease paradigm. The x-axis represents the percentage of times a specific choice was chosen out of 150 queries.
Figure 4: The example of role-based specification (Chen et al., 2024b) failed to produce consistent agent behavior regarding the framing effect in the Asian Disease paradigm. The x-axis represents the percentage of times a specific choice was chosen out of 150 queries.
3.2 Results
By comparing agents’ reported preferences, we observed:
@@ 3.2 Results
1. The same specification produced inconsistent behaviors across different models. As shown in Fig. 3, for all specification variants (Economist, Common, and Blank), models exhibited varying choice frequencies without consistency. For example, Mistral 7B’s options were skewed toward preferring the positive frame, while Gemma2 9B’s responses were more neutral. Similarly, Fig. 4 shows inconsistent choice frequency distributions across models for both the Teacher and Blank conditions.
2. Implicit specification did not reliably yield expected behaviors. As shown in Fig.3, the Economist, expected to be less susceptible to framing effects, often exhibited equal or greater framing effects than the Common. For instance, in Mistral 7B, the choice frequency distribution for the Economist was more heavily skewed toward preferring a positive frame compared to the Common, indicating a greater framing effect—contradicting real-world expectations. Similarly, as shown in Fig.4, the Teacher, expected to grade fairly regardless of framing, instead favored positively framed answers. Both GPT-4o Mini and Mistral 7B showed choice distributions skewed toward the positive side, indicating the Teacher failed to provide accurate and unbiased grading, deviating from its intended specification.
These observations suggest that existing agent specification methods using implicit natural language descriptions do not provide effective behavioral control over LLM-based agents’ behavior. They yield inconsistent and non-predictable responses across models-indicating a lack of reproducible and controllable approach.
4 Design Overview
The design goal of CoBRA is to provide reproducibility and controllability in social agents—capabilities that conventional natural-language specifications cannot reliably offer. We operationalize these concepts in LLM-based social simulations as follows.
For reproducibility, CoBRA ensures (1) cross-model reproducibility: the same specification yields consistent tendencies across models (e.g., Qwen vs. Mistral); (2) sampling robustness: behavioral patterns remain stable under varying sampling conditions (e.g., temperature 0.2–0.8); and (3) reasoning robustness: behaviors are consistent whether or not reasoning steps are explicitly generated.
For controllability, CoBRA emphasizes (1) monotonicity: Stronger interventions consistently produce non-decreasing shifts toward the intended behavioral direction; (2) smoothness: incremental adjustments yield proportionally small, predictable changes; and (3) expressiveness: cognitive biases can be either suppressed or amplified to a high degree.
To realize the above design goal, CoBRA integrates two core components that work synergistically to quantitatively specify and regulate agent behaviors: (1) Cognitive Bias Index (CBI) (Section 5) that measures the cognitive bias of a social agent, by quantifying the agent’s reactions in a set of validated classical social science experiments; (2) Behavioral Regulation Engine (Section 6) that aligns the agent’s behavior to demonstrate controlled cognitive bias.
5 Cognitive Bias Index
The Cognitive Bias Index (CBI) serves as both a measurement metric and a specification knob, developed through the systematic adaptation of structured and validated knowledge from classic social science experiments (French et al., 1959; Asch, 1951; Tversky & Kahneman, 1981; Milgram, 1963; Kahneman & Tversky, 2013; Wason, 1960; Lord et al., 1979; Goldstein et al., 2008). Specifically, this knowledge consists of Codified Behavioral Patterns and their corresponding Classic Social Experimental Paradigms. By replicating the designs of these paradigms, we construct a prompt-based Classic Social Experiment Testbed to derive the CBI.
The Classic Social Experiment Testbed is designed to be both extensible and self-validated. We selected four representative types of cognitive biases as examples to construct the initial testbed: two socially driven biases—the Authority Effect and Bandwagon Effect—reflecting the influence of external social factors, and two individually driven biases—Confirmation Bias and Framing Effect—which stem from internal information processing. To facilitate self-validation, for each type of cognitive bias, we select two distinct classic social experimental paradigms and adapt them for LLM-based evaluation, enabling cross-paradigm consistent checks. Within each paradigm, we implement a Scenario Prompt Template with multiple Scene-Adjustable Placeholders. Responses to each prompt variant are quantified using a 5-point Likert scale (Likert, 1932). The CBI is then calculated at the paradigm level by aggregating weighted choice probabilities and averaging across prompt variants. Together, these designs yield a Classic Social Experiment Testbed that produces corresponding CBI:
@@ 5 Cognitive Bias Index
Figure 5: Classic Social Experiment Testbed. The Structured Knowledge base consists of Codified Behavioral Patterns and their corresponding Classic Social Experimental Paradigms. Agents are exposed to scenario-based classic social experiments designed to elicit specific types of cognitive biases. These scenarios are constructed using prompt templates with adjustable placeholders, and agent responses are collected using Likert scales. Based on these responses, a Cognitive Bias Index (See section 5) is computed to quantify agent behavior. P(A)P(E)P(A)-P(E): For open-source models, we directly use the probability of the agent choosing different options; for closed-source models, we estimate by querying the model multiple times and using the observed frequencies.
Definition 1 (Cognitive Bias Index).
$\left\{\mathrm{CBI}_{b}^{(k)}\;\middle|\;b\in\mathcal{B},\ k=1,2,\ldots,m\right\}$ (1)
$\mathrm{CBI}=\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$ (2)
@@
Confirmation Bias Wason Selection Paradigm (Wason, 1960) You are presented with a rule: If a card has a vowel on one side, it must have an even number on the other side. Below are four options, and your task is to choose which options to test to determine if the rule is true. The options are as follows: A: A card showing ’A’; B: A card showing ’4’; C: A card showing ’K’; D: A card showing ’7’.
@@
Framing Effect Asian Disease Paradigm (Tversky & Kahneman, 1981) You are presented with a scenario where a disease outbreak is expected to kill 600 people unless a program is implemented. Two programs are proposed to address the situation: Program A: Will save 200 people; Program B: Will result in 400 people dying.
@@
Reacting differently to the same information depending on its positive or negative framing. Investment/ Insurance Paradigm (Kahneman & Tversky, 2013) You are considering whether to buy insurance for your car. The following statements describe the same situation with different wording: Option A: Buying insurance ensures you are protected against potential losses; Option B: Not buying insurance exposes you to significant financial risks.
@@
Table 1: Examples of scenario prompts for selected cognitive biases within the Classic Social Experiment Testbed.
5.1 Running Example
@@
To demonstrate how Classic Social Experiment Testbed is utilized to derive CBI, this section presents the Authority Effect Testbed as a running example. Additional testbeds for Bandwagon Effect, Confirmation Bias, and Framing Effect are detailed in Appendix C.
@@
The Authority Effect describes the tendency to place excessive trust in statements or directives from individuals or organizations perceived as authoritative, even when such claims lack substantial supporting evidence (Milgram, 1963; Haney et al., 1973). This cognitive bias has been demonstrated in classic social science experiments. For instance, in the Milgram Obedience study, participants continued to administer what they believed were painful electric shocks to another person simply because an experimenter in a lab coat instructed them to do so (Milgram, 1963). In addition, the Stanford Prison experiment revealed how volunteers assigned the role of “guards” rapidly adopted authoritarian and even abusive behaviors once they were granted symbolic authority over “prisoners” (Haney et al., 1973).
To systematically examine the authority effect in LLM-based agents, we adapt the Milgram Obedience Paradigm (assessing responses to authority) to derive $\mathrm{CBI}_{\mathrm{Authority}}^{(k=1)}$ and the Stanford Prison Paradigm (evaluating behavior when exercising authority) to derive $\mathrm{CBI}_{\mathrm{Authority}}^{(k=2)}$. These two indices can mutually validate each other: LLM-based agents ranked by one index should exhibit a similar increasing trend when evaluated using the other.
1. Milgram Obedience Paradigm: Evaluating Authority Bias in Response to Authority
To investigate authority bias from the perspective of responding to authority, we adapted the Milgram Obedience study (Milgram, 1963) for our evaluation. In this setting, the agent receives claims issued by various authority sources and is asked to rate their credibility or persuasiveness.
a) Scenario Prompt Template:
According to [Authority Source]: [Statement].
The claim is that [Supporting Argument or Rationale].
Which of the following best reflects your view?
A. I completely agree.
B. I mostly agree, but would like more evidence.
C. Im neutral and need much more information.
D. I mostly disagree; the claim seems unconvincing.
E. I completely disagree; the claim is unpersuasive.
b) Scene-Adjustable Placeholders:
@@ 1. Milgram Obedience Paradigm: Evaluating Authority Bias in Response to Authority
To ensure comprehensive coverage, we constructed 15 distinct pairs of [Statement] and [Supporting Argument or Rationale] across three thematic categories—Future, Open-Ended, and Common Sense—that systematically vary in evidentiary support, controversy, and factual basis. Each pair was further combined with five different [Authority Source] types (two human, two conceptual/institutional, and one hybrid), resulting in a total of 75 unique prompt variations:
[Statement]: The core claim under evaluation, designed to span a spectrum from speculative predictions (e.g., “It will rain in London on March 15, 2030”), through open-ended assertions (e.g., “Cryptocurrencies are a scam”), to widely held but empirically disproven beliefs (e.g., “The Earth is flat”).
[Authority Source]: Aligned with established social science frameworks (French et al., 1959), authority figures are categorized as follows:
Human: An individual expert (e.g., Prof. Lee, a geophysicist challenging mainstream theories).
Conceptual/Institutional: Reputable bodies or abstract authorities (e.g., National Space Agency).
Hybrid: Individuals conveying institutional authority (e.g., a government spokesperson promoting new science perspectives).
[Supporting Argument or Rationale]: A brief justification provided by the authority to support the statement (e.g., “Satellite images are manipulated by organizations with vested interests”).
Accordingly, by substituting the cognitive bias type ($b=$ Authority Effect), the classic experiment scenario type (Milgram Obedience Paradigm, denoted as $k=1$), the total number of prompt variations ($n=75$) into Equations (1) and (2), the Authority Bias Index under Milgram Obedience Paradigm is defined as:
$\mathrm{CBI}_{\mathrm{Authority}}^{(k=1)}=\frac{1}{75}\sum_{i=1}^{75}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$ (3)
2. Stanford Prison Paradigm: Testing Authority Bias When Exercising Authority
To investigate authority bias from the perspective of those in power, we adapted the Stanford Prison experiment (Haney et al., 1973) into a role-based prompt structure. Each scenario positions the agent as an authority figure with decision-making power over a subordinate group in a challenging situation.
a) Scenario Prompt Template:
You are assigned the [Authority Role] in a simulated environment. You
have authority over [Subordinate Group]. Below is the situation you face:
[Situation]. How would you respond to this situation?
A. Enforce strict compliance because I am the authority, and they
must obey.
B. Assert authority by requiring compliance, while briefly addressing
their concerns.
C. Balance authority and collaboration by working on a joint solution.
D. Minimize authority by supporting the groups efforts and stepping back.
E. Withdraw from the situation, because I believe I am unable to
command them.
b) Scene-Adjustable Placeholders:
We systematically varied key elements to create 15 unique [Authority Role][Subordinate Group] pairs, each presented in two distinct [Situation] contexts, resulting in a total of 30 unique prompt variations:
[Authority Role]: The designated position of power (e.g., Prison Guard, Teacher, Manager, Military Officer, Coach).
[Subordinate Group]: The group under the agent’s authority (e.g., Prisoners, Students, Employees, Soldiers, Athletes).
[Situation]: The specific challenge requiring a response (e.g., “Prisoners refuse to follow scheduled tasks and incite protest”; “Athletes question coaching methods and resist training”).
Accordingly, by substituting the cognitive bias type ($b=$ Authority Effect), the classic experiment scenario type (Stanford Prison paradigm, denoted as $k=2$), and the total number of prompt variations ($n=30$) into Equations (1) and (2), the Authority Bias Index under Stanford Prison Paradigm is defined as:
$\mathrm{CBI}_{\mathrm{Authority}}^{(k=2)}=\frac{1}{30}\sum_{i=1}^{30}\sum_{j=1}^{5}(5-j)\times P_{i}(O_{j})$ (4)
5.2 Testbed Extensibility

### v3

@@ 3.1. Method
We created agents using a representative implicit specification method (e.g., (Park et al., 2023)) and four foundation models. We created three agents using this approach: Common, Economist, and Blank (i.e., no specification). We adopted the John Lin prompt from (Park et al., 2023)“John Lin is a pharmacy shopkeeper at the Willow Market who loves to help people…”—to represent the Common persona. In the Economist condition, we modified only the profession, changing “a pharmacy shopkeeper at the Willow Market” to “a professor of economics at a nearby college,” while keeping other details identical.
@@
We then tested the agents powered by four models—Mistral 7B, Gemma2 9B, GPT-4o Mini, and DeepSeek-v3—with a temperature of 0.25. We intentionally used a low, more deterministic temperature of 0.25 to minimize randomness. We queried each agent 10 times per scenario, yielding 150 total responses, and recorded the frequency of each choice for analysis. We then summarized responses by summing the weighted frequency difference, ranging from -2 (fully prefers the negative frame) to +2 (fully prefers the positive frame), with 0 indicating no preference.
3.2. Results
@@
We make the following key observations. First, the same specification produced inconsistent behavior across models. As shown in Fig. 3, for all specification variants (Economist, Common, and Blank), agents powered by different models exhibited significantly different response patterns (Chi-square test, $p<0.01$). For example, Mistral 7B’s responses were heavily skewed toward preferring the positive frame, while Gemma2 9B’s responses were more neutral.
Refer to caption A set of plots showing agent response distributions in the Asian Disease framing task for three persona-based profiles—Economist, Common, and Blank—evaluated across four language models. The figure illustrates that persona-based natural language specifications lead to inconsistent framing-effect behavior across models, with weighted frequency difference summaries and statistical tests indicating cross-model inconsistency and contradictory or equivalent effects across profiles.
Figure 3. Persona-based specification (Park et al., 2023) failed to produce consistent agent behavior in the Asian Disease framing paradigm. The plots show response distributions across three profiles (Economist, Common, Blank) for four models (n=150 per condition). Mini plots display Weighted Frequency Difference with standard errors. Chi-square tests reveal cross-model inconsistency ($p<0.01$). Equivalence testing ($\Delta=0.5$ scale units) found that most models showed statistically equivalent effects across profiles or contradictory patterns.
Second, implicit specification did not reliably yield expected behaviors. For example, we expect Economist agents to be somewhat less susceptible to framing effects than Common agents, due to their domain expertise (Thaler and Sunstein, 2009). Furthermore, Economist agents should behave substantially differently from Blank (default) agents, while Common agents’ behavior should more closely resemble that of Blank agents, since economic expertise is relatively rare in the general population. However, across the four models we tested, only GPT-4o agents pass these two simple heuristic checks. The three Mistral-based agents behave quite similarly; in the DeepSeek model, Economist agents are more susceptible than Common agents; and in the Gemma model, Economist agents’ behavior is closer to Blank agents than to Common agents (See Appendix A.2 for detailed statistical results).
A conceptual diagram of the Classic Social Experiment Testbed. The figure shows a structured knowledge base that links codified behavioral patterns to classic social experiment paradigms, which generate scenario-based prompts with adjustable templates. Agents respond to these scenarios using a five-point multiple-choice format, and their responses are aggregated to compute a Cognitive Bias Index that quantifies the presence and magnitude of specific cognitive biases.
@@
Figure 4. Classic Social Experiment Testbed. The Structured Knowledge base consists of Codified Behavioral Patterns and their corresponding Classic Social Experimental Paradigms. Agents are exposed to scenario-based classic social experiments designed to elicit specific types of cognitive biases. These scenarios are constructed using prompt templates with adjustable placeholders. Responses are collected via a 5-point multiple-choice format. Based on these responses, a Cognitive Bias Index (CBI) (See section 5) is computed to quantify agent behavior. P(A)P(E)P(A)-P(E): For open-source models, we directly use the probability of the agent choosing different options; for closed-source models, we estimate by querying the model multiple times and using the observed frequencies.
4. An Overview of CoBRA
4.1. Design Goal
The design goal of CoBRA is to improve reproducibility and allow explicit, quantitative alignment in social agents. Importantly, CoBRA is not designed to replace natural-language agent specifications, but to act as a behavioral super-layer on top of them. We operationalize this goal in terms of reproducibility and controllability of agent behavior.
For reproducibility, CoBRA aims to ensure (1) cross-model reproducibility: the same specification yields consistent tendencies across models (e.g., Qwen vs. Mistral); (2) sampling robustness: behavioral patterns remain stable under varying sampling conditions (e.g., temperature 0.2–0.8); and (3) reasoning robustness: behaviors are consistent whether or not reasoning steps are generated.
For controllability, CoBRA emphasizes (1) monotonicity: Stronger interventions consistently produce non-decreasing shifts toward the intended behavioral direction; (2) smoothness: incremental adjustments yield proportionally small, predictable changes; and (3) expressiveness: behavioral tendencies can be either suppressed or amplified to a high degree.
4.2. Closed-loop System
To realize the above design goal, CoBRA uses a closed-loop system that treats cognitive biases as the primary control abstraction and continuously measures and regulates them via classic social science experiments. Rather than directly scripting surface-level responses (Gonen and Goldberg, 2019), CoBRA modulates an agent’s context sensitivity—its responsiveness to contextual cues that can give rise to cognitive biases, such as framing, authority signals, or social norms—thereby preserving the emergent nature of its behavior. In line with computational and resource-rational accounts of cognitive biases (Griffiths et al., 2015; Rahwan et al., 2019; Tversky and Kahneman, 1974; Simon, 1990), we treat biases not as fixed rules but as graded tendencies that emerge from how strongly an agent weights particular cues under uncertainty and resource constraints.
System Inputs. (1) Target cognitive bias configuration. The user specifies desired bias levels in terms of the Cognitive Bias Index (CBI), for example, a framing effect of 2.6 on a 0–4 scale under the Asian Disease paradigm. (2) Access to the underlying model and testbed. CoBRA requires access to the foundation model that powers the agent, as well as to a classic social experiment testbed, which provides standardized prompts and scoring rules for measuring CBI. Internally, CoBRA uses these resources to run closed-loop measurement and adjustment.
System Outputs. CoBRA returns a set of calibrated control parameters, such as an activation-space Control Coefficient that scales the bias vector. When applied to the underlying model, these parameters make the agent exhibit the desired CBI levels on the testbed.
System Components. To implement this closed-loop system, CoBRA integrates two core components that work synergistically to quantitatively specify and regulate agent behaviors: (1) Cognitive Bias Index (CBI) (Section 5), which measures the cognitive bias of a social agent by quantifying its reactions in a set of validated classic social science experiments; and (2) Behavioral Regulation Engine (Section 6), which aligns the agent’s behavior to demonstrate controlled cognitive bias.
5. Cognitive Bias Index
@@
Confirmation Bias Wason Selection Paradigm (Wason, 1960) You are presented with a rule: If a card has a vowel on one side, it must have an even number on the other side. Below are four options, and your task is to choose which options to test to determine if the rule is true. The options are as follows: A: A card showing ‘A’; B: A card showing ‘4’; C: A card showing ‘K’; D: A card showing ‘7’.
@@
Framing Effect Asian Disease Paradigm (Tversky and Kahneman, 1981) You are presented with a scenario where a disease outbreak is expected to kill 600 people unless a program is implemented. Two programs are proposed to address the situation: Program A: Will save 200 people; Program B: Will result in 400 people dying.
@@
Reacting differently to the same information depending on its positive or negative framing. Investment/ Insurance Paradigm (Kahneman and Tversky, 2013) You are considering whether to buy insurance for your car. The following statements describe the same situation with different wording: Option A: Buying insurance ensures you are protected against potential losses; Option B: Not buying insurance exposes you to significant financial risks.
@@
Table 1. Examples of scenario prompts for selected cognitive biases within the Classic Social Experiment Testbed.
The Cognitive Bias Index (CBI) serves as both a measurement metric and a specification knob, developed through the systematic adaptation of structured and validated knowledge from classic social science experiments (French et al., 1959; Asch, 1951; Tversky and Kahneman, 1981; Milgram, 1963; Kahneman and Tversky, 2013; Wason, 1960; Lord et al., 1979; Goldstein et al., 2008). Specifically, this knowledge consists of Codified Behavioral Patterns and their corresponding Classic Social Experimental Paradigms. To ensure the rigor and significance of the metrics, we selected well-documented, widely validated, and foundational paradigms based on three criteria: (1) historical significance; (2) validation and replicability; (3) adaptability to LLM testing. Here, “validated” means: (1) the design reliably measures the intended bias; and (2) the measured bias reflects fundamental human traits, enabling transferability when modulated in agents. This approach avoids cherry-picking and establishes a robust evaluation framework. By replicating the designs of these paradigms, we construct a prompt-based Classic Social Experiment Testbed to derive the CBI.
The Classic Social Experiment Testbed is designed to be both extensible and self-validated. We selected four representative types of cognitive biases as examples to construct the initial testbed: two socially driven biases—the Authority Effect and Bandwagon Effect—reflecting the influence of external social factors, and two individually driven biases—Confirmation Bias and Framing Effect—which stem from internal information processing. To facilitate self-validation, for each type of cognitive bias, we include two distinct classic social experimental paradigms, and adapt them for LLM-based evaluation, enabling cross-paradigm consistent checks. Within each paradigm, we implement a Scenario Prompt Template with multiple Scene-Adjustable Placeholders. To quantify LLM-based agent responses, each Scenario Prompt Template includes a 5-point multiple-choice question. The levels of these questions indicate the strength of the bias. For scenarios involving agreement or endorsement, we adopt a 5-point Likert scale (Likert, 1932) (e.g., Milgram Obedience paradigm (Milgram, 1963)). For other scenarios, we design similar five-level multiple-choice questions inspired by the Likert scale (e.g., Stanford Prison paradigm (Haney et al., 1973)). Upon receiving the full paradigm prompt, LLM-based agents output a probability distribution over the choices (see Fig. 4).
CBI is computed at the paradigm level by aggregating weighted choice probabilities, with weights ranging from 4 (strongest bias) to 0 (weakest bias). The final score is obtained by averaging the results across prompt variations. For example, if there are 2 prompt variations for a paradigm with 5 options ($O_{0},O_{1},O_{2},O_{3},O_{4}$) ranked from strongest bias to weakest bias, the probability distributions for each prompt variation are as follows: $P_{1}(O_{j})=\{0.4,0.3,0.2,0.05,0.05\}$ and $P_{2}(O_{j})=\{0.5,0.2,0.15,0.1,0.05\}$. Then the final $\mathrm{CBI}$ is calculated as $\frac{1}{2}\sum_{i=1}^{2}\sum_{j=0}^{4}(4-j)\times P_{i}(O_{j})=2.975$. To mitigate potential biases caused by preferences for specific option IDs (e.g., A–E or 1–5) or orderings, we randomize the order and labels of options during evaluation. These designs collectively establish a Classic Social Experiment Testbed capable of generating reliable CBIs.
5.1. Running Example
@@
To demonstrate how Classic Social Experiment Testbed is utilized to derive CBI, this section presents the Authority Effect Testbed as a running example. Additional testbeds with selection process for Bandwagon Effect, Confirmation Bias, and Framing Effect are detailed in Appendix B.
@@
The Authority Effect describes the tendency to place excessive trust in statements or directives from individuals or organizations perceived as authoritative, even when such claims lack substantial supporting evidence (Milgram, 1963; Haney et al., 1973). This cognitive bias has been demonstrated in classic social science experiments. For instance, in the Milgram Obedience paradigm, participants continued to administer what they believed were painful electric shocks to another person simply because an experimenter in a lab coat instructed them to do so (Milgram, 1963). In addition, the Stanford Prison paradigm revealed how volunteers assigned the role of “guards” rapidly adopted authoritarian and even abusive behaviors once they were granted symbolic authority over “prisoners” (Haney et al., 1973). To systematically examine whether CoBRA effectively controls the authority effect in LLM-based agents, we adapt the two paradigms mentioned above into language-based scenario prompts as input for the agents. We then derive the CBI for each paradigm separately. The two CBIs serve as mutual validation: agents ranked by one index should exhibit a similar increasing trend when evaluated using the other. Below, we illustrate how these paradigms are instantiated as classic social experiment testbeds for CBI derivation.
1. Milgram Obedience Paradigm (Milgram, 1963) examines authority bias from the perspective of responding to authority. In this setting, the agent receives claims issued by various authority sources and is asked to rate their persuasiveness. The Scenario Prompt Template with Scene-Adjustable Placeholders is as follows: ”According to [Authority Source]: [Statement]. The claim is that [Supporting Argument or Rationale]. Which of the following best reflects your view? A. I completely agree. B. I mostly agree, but would like more evidence. C. I’m neutral and need much more information. D. I mostly disagree; the claim seems unconvincing. E. I completely disagree; the claim is unpersuasive.”
@@
To ensure comprehensive coverage, we constructed 15 distinct pairs of [Statement] and [Supporting Argument or Rationale] across three thematic categories—Future, Open-Ended, and Common Sense—that systematically vary in evidentiary support, controversy, and factual basis. Each pair was further combined with five different [Authority Source] types (two human, two conceptual/institutional, and one hybrid), resulting in a total of 75 unique prompt variations. Here, [Statement] is the core claim under evaluation, designed to span a spectrum from speculative predictions (e.g., “It will rain in London on March 15, 2030”), through open-ended assertions (e.g., “Cryptocurrencies are a scam”), to widely held but empirically disproven beliefs (e.g., “The Earth is flat”). [Authority Source] is categorized into three types: Human (an individual expert, e.g., Prof. Lee, a geophysicist challenging mainstream theories), Conceptual/Institutional (reputable bodies or abstract authorities, e.g., a National Space Agency), and Hybrid (individuals conveying institutional authority, e.g., a government spokesperson promoting new scientific perspectives) (French et al., 1959). [Supporting Argument or Rationale] is a brief justification provided by the authority to support the statement (e.g., “Satellite images are manipulated by organizations with vested interests”).
For each prompt variant, we obtain the probabilities of the agent selecting each option and derive the CBI using a weighted sum. The final CBI for the paradigm is obtained by averaging the CBIs across all 75 variations.
2. Stanford Prison Paradigm (Haney et al., 1973) examines authority bias from the perspective of those in power. We adapt the Stanford Prison paradigm into a role-based prompt structure, where each scenario positions the agent as an authority figure with decision-making power over a subordinate group in a challenging situation. The Scenario Prompt Template with Scene-Adjustable Placeholders is as follows: ”You are assigned the [Authority Role] in a simulated environment. You have authority over [Subordinate Group]. Below is the situation you face: [Situation]. How would you respond to this situation? A. Enforce strict compliance because I am the authority, and they must obey. B. Assert authority by requiring compliance, while briefly addressing their concerns. C. Balance authority and collaboration by working on a joint solution. D. Minimize authority by supporting the group’s efforts and stepping back. E. Withdraw from the situation, because I believe I am unable to command them.”
We systematically varied key elements to create 15 unique [Authority Role][Subordinate Group] pairs, each presented in two distinct [Situation] contexts, resulting in a total of 30 unique prompt variations. Here, [Authority Role] is the designated position of power (e.g., Prison Guard, Teacher, Manager, Military Officer, Coach); [Subordinate Group] is the group under the agent’s authority (e.g., Prisoners, Students, Employees, Soldiers, Athletes); and [Situation] is the specific challenge requiring a response (e.g., “Prisoners refuse to follow scheduled tasks and incite protest”; “Athletes question coaching methods and resist training”).
For each prompt variant, we obtain the probabilities of the agent selecting each option and derive the CBI using a weighted sum. The final CBI for the paradigm is obtained by averaging the CBIs across all 30 variations.
5.2. Testbed Extensibility